OpenStudy (anonymous):

i will fan and medal plz help 4.8+x-5.9=3.6

OpenStudy (anonymous):

collect like terms 4.8-5.9+x=3.6

OpenStudy (anonymous):

x-1.1=3.6 x=3.6+1.1 =4.7

OpenStudy (anonymous):

ty so much

OpenStudy (anonymous):

solve 4.8 + x - 5.9 = 3.6 ⇒ x=4.7

OpenStudy (anonymous):

it was right ty